OpenStudy (anonymous):

Convert to a fraction: 75%

OpenStudy (anonymous):

75% = 75/100 = 3/4

OpenStudy (anonymous):

you can do that with any number. I think i understand fractions better now. thanks hawk.

OpenStudy (anonymous):

sure thing :)

OpenStudy (dmezzullo):

75% = 3/4

OpenStudy (anonymous):

percent means 1/100 therefore, 75% means 75*1/100 which equals 75/100. if you reduce this equation by dividing by 5 you get 3/4